# SPADE: Spatial Transcriptomics and Pathology Alignment Using a Mixture of Data Experts for an Expressive Latent Space

### v1

@@ 4.2 Baselines and Ablations
We compared our approach against three baseline encoders: UNI, TANGLE (pancancer-trained) (Section 2), CONCH [18] and Prov-GigaPath [32]. We assessed two distinct aggregation mechanisms for slide-level features: ABMIL and TransMIL.
@@
Alongside the baseline comparisons, we conducted an ablation study to compare methods for routing data experts (Section 3.4). Additionally, we conducted an ablation study to experiment with different numbers of clusters in both the initial and second steps of clustering. Lastly, we implemented a version of SPADE with one-step K-means clustering to evaluate the effect of the additional clustering step on downstream performance (see Supplementary Materials).
To thoroughly evaluate SPADE, we conducted a comparison with a single-expert model (SPADE-0) across subtyping, survival, and BCR prediction tasks (see Supplementary Materials).
@@ 4.3 Implementation Details
For the classification ABMIL and TransMIL models, we trained them from scratch using the same set of hyperparameters across all downstream experiments (see Supplementary Materials)
@@ 4.4 Results
Prostate BCR Lung Uterine Breast
Private TCGA-PRAD PLCO TCGA-UCEC TCGA-BRCA PLCO
c-index c-index c-index c-index c-index c-index
$p_{1}$+ABMIL 63.7 57.6 56.6 70.7 68.3 56.6
$p_{2}$+ABMIL 66.8 59.0 57.8 71.5 69.5 57.0
$p_{3}$+ABMIL 67.0 59.4 64.7 72.4 71.8 59.5
$p_{1}$+TransMIL 66.2 57.9 67.7 70.2 69.4 517
$p_{2}$+TransMIL 69.7 58.1 71.7 70.4 71.6 52.9
$p_{3}$+TransMIL 71.4 58.8 72.3 71.0 71.8 56.9
@@ 4.4.1 Cancer Subtyping
The SPADE framework consistently outperforms baseline models across a range of cancer subtyping tasks, showing clear advantages in both AUROC and F1-score metrics. Notably, SPADE achieves the highest AUC across six out of eight tasks and secures the highest F1 score for six of the tasks (Table 1). CONCH features led to higher AUC for lung (CPTAC) and breast (TCGA-BRCA) subtyping tasks, as well as higher F1 score for lung subtyping (PLCO). UNI features resulted in a higher F1 score for the ovarian cancer suptyping task (UBC-OCEAN). The comparison with TANGLE’s performance highlights the distinct advantage of spatial transcriptomics over Bulk RNA-seq. Using Bulk RNA-seq may limit model expressibility, likely due to the loss of heterogeneity and the lack of one-to-one alignment between expression and imaging data.
Lung Prostate Lymph Nodes Ovarian Breast
PLCO CPTAC PANDA
TCGA-
PRAD
CAM16 UBC-OCEAN PLCO
TCGA-
BRCA
AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC F1
[…]
$16(k_{1}),16(k_{2})$,TransMIL 96.5 82.0 96.3 89.7 89.5 58.9 72.0 43.9 93.4 82.9 97.3 83.5 72.5 47.9 97.3 98.7
$16(k_{1}),8(k_{2})$,TransMIL 94.1 83.7 93.5 85.3 89.9 54.6 78.1 30.5 91.3 80.3 97.4 83.1 78.0 55.4 97.0 96.8
[…]
Prostate BCR Lung Uterine Breast
Private
TCGA-PRAD
PLCO
TCGA-UCEC
TCGA-BRCA
PLCO
c-index c-index c-index c-index c-index c-index
[…]
$16(k_{1}),8(k_{2})$,ABMIL 66.2 55.7 53.2 70.4 70.8 56.1
$16(k_{1}),32(k_{2})$,ABMIL 70.2 57.4 60.7 71.7 71.7 57.0
[…]
$16(k_{1}),8(k_{2})$,TransMIL 65.3 52.9 61.4 68.9 70.7 52.9
$16(k_{1}),32(k_{2})$,TransMIL 65.8 52.5 62.4 70.6 71.1 52.5
[…]
4.4.2 Survival Prediction and Prostate BCR Analysis
SPADE outperforms all baseline models, achieving the highest c-index for three out of the four survival tasks and all BCR prediction tasks. In the TCGA-BRCA survival task, UNI surpasses SPADE, but SPADE still achieves the second-highest performance. SPADE outperforms TANGLE on all datasets, demonstrating the critical importance of spatial context in modeling disease progression, which is achieved by incorporating ST data.
4.4.3 Ablation Studies
For the data experts’ routing ablation studies, there is varying performance between different routing techniques (Table 3, 4). For most subtyping datasets, $p_{2}$ (averaging) and $p_{3}$ (distance weighting) routing methods perform better than $p_{1}$, which simply uses the expert of the closest centroid. This indicates the additional model input from other experts provides a performance advantage, perhaps allowing for more complex representations of the WSI.
Our ablation study on the two-step clustering procedure revealed that varying the number of clusters at each stage ($k_{1}$ and $k_{2}$) can influence the model performance, but the improvement varies across different datasets and cancer types (see Table 5, 6). For the subtyping task, the configuration with $k_{1}=16,k_{2}=16$ using ABMIL achieved the highest AUC in five datasets, demonstrating the effectiveness of balanced hierarchical clustering. Increasing the second-stage clustering granularity ($k_{2}=32$) improved F1 scores in several cases, such as lung PLCO, ovarian, and breast TCGA-BRCA, suggesting that a finer-grained latent space can enhance classification precision, but it also requires double the computation resources or double the runtime. Reducing the second-stage clustering granularity ($k_{2}=8$) led to mixed results: while it improved F1 in some datasets, it generally resulted in lower AUC scores.
For the survival and BCR, $k_{1}=16,k_{2}=16$ achieved the highest c-index with ABMIL and TransMIL, demonstrating the effectiveness of balanced clustering. Increasing $k_{1}$ and $k_{2}$ to 32 improved the c-index for TCGA-BRCA while reducing $k_{2}$ to 8 led to performance drops across all survival tasks.
Further comparisons between SPADE and its single-expert version highlight the significant impact of the mixture of data experts module on the performance, as evidenced by SPADE’s consistently higher AUC and F1 score (see Supplementary Materials). This trend may be attributed to the diverse tissue types represented in the HEST-1k dataset (e.g. brain, heart, lung). A model trained on all organ types might lack the specificity needed to improve performance for each task. Given the high variability of genes present between tissue types, training on data from all organs within a single model may limit its ability to capture the heterogeneity essential for effective subtyping and survival prediction. Additionally, as discussed in Section 2.2, with such diverse tissue types, contrastive learning could rely on easy negative pairings to drive down the loss, resulting in a less expressive latent space.
4.4.4 Interpretability Analysis
We conducted interpretability analyses by plotting the attention weights from the ABMIL variants on our in-house Prostate BCR dataset to better understand the differences between the SPADE and UNI features. In Figure 2, we present attention heatmaps, where deeper red indicates higher attention weights, and the predicted cancer mask is shown in yellow. We validated this cancer prediction mask against pathology reports. In panel (a), UNI attention highlights some small regions within the tumor boundary, but much of the attention is diffuse. In contrast, the SPADE-N framework’s attention weights display higher attention regions within the cancer region and a hot zone along the bottom right border of the tumor mask.
The heatmap 2 depicts what regions of the WSI had the highest attention for a case from the in-house BCR task. The attention heatmap produced by SPADE highlights areas predominantly showing poorly formed and compact $4+5$ and $4+3$ tumor patterns with tertiary $5$ according to the Gleason grading system [7], subtle desmoplastic stromal response with slightly cellular fibroblasts exhibiting pale cytoplasm and nuclear enlargement.
5 Conclusions
Based on the performance across our 14 downstream tasks, the SPADE framework has shown to be an expressive feature encoder using spatial transcriptomics data. Utilization of HEST-1k allowed our model to learn robust, biologically relevant embeddings by aligning histological image features with spatially resolved gene expression patterns, which set a strong foundation for downstream pathology analysis tasks. Our improved performance over the single expert version of our framework highlights the benefit of SPADE’s mixture of experts approach. Additionally, our results show significant improvement over TANGLE, which relies on bulk sequencing data, underscoring the value of spatial gene profiling. SPADE’s ability to utilize spatial context allows our model to capture localized differences in gene expression that are crucial for accurate disease subtyping and survival prediction.
Supplementary Material
S6 Datasets
Prostate For our prostate cancer biochemical recurrence (BCR) tasks, we use an in-house dataset of radical prostatectomy whole-mount WSIs and the TCGA-PRAD dataset. The in-house dataset spans 161 patients, with one whole-mount WSI selected per patient based on pathology reports. The ratio of patients without BCR versus those with BCR is 135:26. The Gleason grading breakdown of this dataset is: 74 - 3+4 cases, 46 - 4+3 cases, 20 - 4+5 cases, 13 - 3+3 cases, 6 - 4+4 cases, 1 - 3+5 case, and 1 - 5+4 case. The TCGA-PRAD dataset consists of 336 WSIs spanning 322 patients with a BCR rate ratio of 285:37. We perform a patient-level, label-stratified split into training, validation, and test sets with a 70:10:20 ratio for both in-house and TCGA-PRAD datasets. Performance was evaluated using the c-index.
S7 Model and training
TransMIL downstream architecture: TransMIL architecture was used in the downstream experiments. The core of the TransMIL architecture features two Transformer layers, both configured with a hidden dimension of 256 or 512. After these Transformer layers, a layer normalization step is applied with a dimensionality of 256 or 512. Finally, the architecture includes a classification layer and a linear transformation mapping from 256 or 512 features to the number of target classes.
S8 Clustering
In the first clustering step, we performed clustering within each tissue type, varying $k$ across a range of values. To determine the optimal number of clusters, we calculated the within-cluster sum of squares (WCSS) and applied the elbow method. The corresponding elbow plot (see Figure S3), suggests that the optimal choice is $k=16$, as improvements in WCSS diminish beyond this point. A secondary drop is observed at $k=32$, which we included in our ablation studies to assess its impact on downstream task performance.
In the second clustering step, we clustered the centroids obtained prior to form broader, overall clusters. Again, we evaluated values of $k$ using WCSS and generated an elbow plot (see Figure S3). In this case, the distinction between $k=16$ and $k=32$ is less pronounced. However, given the higher computational cost of running 32 experts, we opted for $k=16$ to reduce overall model complexity. This approach allowed us to structure the data efficiently while balancing computational cost and cluster granularity.
Figure S3: Within-Cluster Sum of Squares (WCSS) Elbow Plots Across Values of K A. First-level clustering performance evaluated using normalized WCSS across 14 organs. B. Second-level clustering performance evaluated using normalized WCSS.
S9 Additional experiments
We observe that data experts trained on cluster centroids obtained with a two-step clustering approach lead to better results compared to a one-step clustering approach (Table S7). The one-step approach results in a reduction in AUC for the CPTAC dataset, with decreases of 1.9% for SPADE+ABMIL and 0.62% for SPADE+TransMIL, and for the Camelyon dataset, with decreases of 5.68% for SPADE+ABMIL and 2.46% for SPADE+TransMIL. F1 values decreased by 3.03% for SPADE+ABMIL and 0.45% for SPADE+TransMIL on the CPTAC dataset. For the Camelyon dataset, F1 values dropped by 12.91% for SPADE+ABMIL and 16.91% for SPADE+TransMIL. This can be explained by the fact that one-step clustering leads to cluster centers that are not sufficiently representative of a specific tissue prototype and, therefore, cannot learn a good representation.
To further evaluate the effectiveness of SPADE, we conducted an experiment comparing its performance against a single-expert model (SPADE-0) across subtyping, survival and BCR prediction tasks. As shown in Tables S8, S9, SPADE consistently outperformed SPADE-0 in most datasets, demonstrating the advantage of incorporating multiple experts tailored to different tissue types.
Lung Lymph Nodes
CPTAC Camelyon
AUC F1 AUC F1
SPADE+ABMIL 96.6 89.7 93.6 77.6
SPADE+TransMIL 96.0 89.3 91.1 68.9
Table S7: One-step clustering for the mixture of experts. We compare the performance of SPADE with 240 data experts using two subtyping tasks. Metrics AUC and F1 are shown as percentages.
@@ 5 Conclusions
Table S8: Comparison of mixture-of-experts (SPADE) to single expert (SPADE-0) for subtyping task. Metrics AUC and F1 are shown as percentages.
Prostate BCR Lung Uterine Breast
Private TCGA-PRAD PLCO TCGA-UCEC TCGA-BRCA PLCO
@@
Table S9: Comparison of mixture-of-experts (SPADE) to single expert (SPADE-0) for survival and BCR. c-index values are shown as percentages.
S10 Limitations
While we trained on the largest and most diverse public Visium dataset available, the sample size 515 remains small compared to large digital pathology datasets. However, given the cost of acquiring spatial transcriptomics data and the relative novelty of the technique, it may be some time before datasets of comparable size to digital pathology collections can be assembled.

### v2

@@ 4.2 Baselines and Ablations
We compared our approach against three baseline vision encoders: UNI, TANGLE (pancancer-trained) (Section 2), CONCH [31] and Prov-GigaPath [8].
@@
Alongside the baseline comparisons, we conducted an ablation study to evaluate different strategies for routing data experts (Sections 3.4, 4.4.4). We further examined the impact of varying the number of clusters in both stages of the two-step clustering procedure (Section 4.4.4), as well as a simplified variant of SPADE using one-step K-means clustering to assess the contribution of the additional clustering step (see E).
To thoroughly evaluate SPADE, we conducted a comparison with a single-expert model (SPADE-0) across subtyping, survival, and BCR prediction tasks (see F).
In addition, we explored the effect of the underlying foundation model by replacing the UNI encoder with the Virchow2 foundation model [32] within SPADE and compared its performance to both standard Virchow2 features and a single-expert version of SPADE trained under the same settings (see G).
@@ 4.3 Implementation Details
We trained the classification ABMIL from scratch using the same set of hyperparameters across all downstream experiments (see B).
@@ 4.4.1 Cancer Subtyping
The SPADE framework consistently outperforms baseline models across a range of cancer subtyping tasks, showing clear advantages in both AUROC and F1-score metrics. Notably, SPADE achieves the highest AUC across seven out of eight tasks and secures the highest F1 score for seven of the eight tasks as well (Table 1). CONCH features led to higher AUC for the breast (TCGA-BRCA) subtyping task. UNI features resulted in a higher F1 score for the ovarian cancer suptyping task (UBC-OCEAN). The comparison with TANGLE’s performance highlights the distinct advantage of spatial transcriptomics over Bulk RNA-seq. Using Bulk RNA-seq may limit model expressibility, likely due to the loss of heterogeneity and the lack of one-to-one alignment between expression and imaging data.
@@
[…]
4.4.2 Survival Prediction and Prostate BCR Analysis
SPADE outperforms all baseline models, achieving the highest c-index for four out of the five survival tasks and the BCR prediction task. In the TCGA-BRCA survival task, UNI surpasses SPADE, but SPADE still achieves the second-highest performance. SPADE outperforms TANGLE on all datasets, demonstrating the critical importance of spatial context in modeling disease progression, which is achieved by incorporating ST data.
Prostate Lung Uterine Breast
UCLA (BCR) TCGA-PRAD PLCO TCGA-UCEC TCGA-BRCA PLCO
@@ 4.4.2 Survival Prediction and Prostate BCR Analysis
UNI+ABMIL $8 \ul5$6.5 57.8 64.6 74.1 56.0
CONCH+ABMIL 60.3 55.9 35.8 64.5 64.5 55.5
[…]
Table 2: Survival and BCR prediction results of SPADE and baselines for six different survival tasks. c-index values are shown as percentages. The best performance is in bold, and the second best is underlined. TANGLE could not be applied to TCGA datasets since they were used to train the model.
4.4.3 Gene Biomarker Prediction
Predicting gene alterations and receptor status from histology slides alone is challenging, and the datasets often suffer from high imbalance. Across the six tasks evaluated, SPADE achieves the highest AUROC on five of the tasks (Table 3). The sole exception is TCGA-CRC BRAF mutation prediction, where CONCH performs slightly better. SPADE also yields the top F1 performance on ER, PR, and TP53 tasks, with CONCH performing better on HER2, GigaPath on KRAS, and TANGLE on BRAF. Averaged over all tasks, SPADE attains the highest mean AUROC (79.4) and mean F1 (64.6).
TCGA-BRCA TCGA-CRC
HER2 ER PR KRAS BRAF TP53
AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC F1
TANGLE - - - - - - - - - - - -
[…]
Table 3: Molecular biomarker prediction results on TCGA cohorts. Comparison of SPADE with baseline methods for predicting receptor status in TCGA-BRCA (HER2, ER, PR) and mutational status in TCGA-CRC (KRAS, BRAF, TP53). Results are reported in terms of AUC and F1 score (%). The best performance is in bold, and the second best is underlined. TANGLE could not be applied to TCGA datasets since they were used to train the model.
4.4.4 Ablation Studies
For the data experts’ routing ablation studies, there is varying performance between different routing techniques (Table 4, 5). For most subtyping datasets, $p_{2}$ (averaging) and $p_{3}$ (distance weighting) routing methods perform better than $p_{1}$, which simply uses the expert of the closest centroid. This suggests that incorporating additional model input from other experts through ensembling provides a performance advantage, potentially enabling more complex representations of the WSI.
Lung Prostate Lymph Nodes Ovarian Breast
PLCO CPTAC PANDA CAM16 UBC-OCEAN PLCO
AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC F1
[…]
$p_{3}$ 98.5 95.5 97.9 92.5 94.9 74.7 98.6 89.1 97.8 87.1 82.1 57.3
Prostate Lung Uterine Breast
UCLA (BCR) TCGA-PRAD PLCO TCGA-UCEC TCGA-BRCA PLCO
c-index c-index c-index c-index c-index c-index
$p_{1}$ 63.7 57.6 56.6 70.7 68.3 56.6
[…]
$p_{3}$ 67.0 59.4 64.7 72.4 71.8 59.5
Our ablation study on the two-step clustering procedure revealed that varying the number of clusters at each stage ($k_{1}$ and $k_{2}$) can influence model performance, but the improvement varies across different datasets and cancer types (see Table 6, 7). For the subtyping task, the configuration with $k_{1}=16,k_{2}=16$ achieved the highest AUC in five datasets, demonstrating the effectiveness of balanced hierarchical clustering. Increasing the second-stage clustering granularity ($k_{2}=32$) improved F1 scores in several cases, such as ovarian, suggesting that a finer-grained latent space can enhance classification precision, but it also requires double the computational resources or double the runtime. Reducing the second-stage clustering granularity ($k_{2}=8$) led to mixed results: while it improved F1 in some datasets, it generally resulted in lower AUC scores.
Lung Prostate Lymph Nodes Ovarian Breast
PLCO CPTAC PANDA
TCGA-
PRAD
CAM16 UBC-OCEAN PLCO
TCGA-
BRCA
AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC F1 AUC F1
[…]
For the survival and BCR, $k_{1}=16,k_{2}=16$ achieved the highest c-index for four out of six tasks, demonstrating the effectiveness of balanced clustering. Increasing $k_{1}$ and $k_{2}$ to 32 improved the c-index for TCGA-BRCA while reducing $k_{2}$ to 8 led to performance drops across all survival tasks.
Prostate Lung Uterine Breast
UCLA (BCR) TCGA-PRAD PLCO TCGA-UCEC TCGA-BRCA PLCO
c-index c-index c-index c-index c-index c-index
[…]
$16(k_{1}),8(k_{2})$ 66.2 55.7 53.2 70.4 70.8 56.1
[…]
Further comparisons between SPADE and its single-expert version highlight the significant impact of the mixture of data experts module on the performance, as evidenced by SPADE’s consistently higher AUC and F1 score (see E). This trend may be attributed to the diverse tissue types represented in the HEST-1k dataset (e.g., brain, heart, lung). A model trained on all organ types might lack the specificity needed to improve performance for each task. Given the high variability of genes present between tissue types, training on data from all organs within a single model may limit its ability to capture the heterogeneity essential for effective subtyping and survival prediction. Additionally, as discussed in Section 2.2, with such diverse tissue types, contrastive learning could rely on easy negative pairings to drive down the loss, resulting in a less expressive latent space.
4.4.5 Interpretability Analysis
We conducted our first interpretability analysis by plotting the attention weights from the ABMIL variants on our in-house UCLA Prostate BCR dataset to better understand the differences between the SPADE and UNI features. In Figure 3, we present attention heatmaps, where deeper red indicates higher attention weights, and the predicted cancer mask is shown in yellow. We validated this cancer prediction mask against pathology reports. In panel (a), UNI attention highlights some small regions within the tumor boundary, but much of the attention is diffuse. In contrast, the SPADE framework’s attention weights display higher attention regions within the cancer region and a hot zone along the bottom right border of the tumor mask.
The heatmap depicts which regions of the WSI had the highest attention for a case from the UCLA BCR task. The attention heatmap produced by SPADE highlights areas predominantly showing poorly formed and compact $4+5$ and $4+3$ tumor patterns with tertiary $5$ according to the Gleason grading system [34], subtle desmoplastic stromal response with slightly cellular fibroblasts exhibiting pale cytoplasm and nuclear enlargement.
Figure 3: Interpretability of SPADE. Attention scores for a prostate WSI positive for biochemical recurrence (BCR) are visualized as a heatmap for the UNI and SPADE models. The deeper the red color, the higher attention the model puts on that region of the tissue. The black outline shows the predicted tumor boundary from our internal cancer prediction model for whole-mount prostectomy WSIs. SPADE shows a higher concentration of attention within the tumor boundary compared to UNI, with some areas of focus along the tumor border. Sub-crops are selected to show the patterns with high attention at higher magnifications.
4.4.6 Comparison of TCGA-CRC Attention Heatmaps from SPADE vs. UNI
To compare SPADE and UNI for gene-related tasks, attention weights from the TCGA-CRC BRAF mutation task were extracted to identify differences in the attentions of each model.
We created a series of heatmaps similar to our setup for Prostate BCR attention visualization (see Figure 4, 7, 8). Additionally, we extracted these heatmaps at high resolution and created a setup for a board-certified pathologist to overlay the heatmap on tissue at full resolution. This setup allowed the pathologist to zoom in on ”hot” regions and examine the underlying tissue patterns. From the heatmaps, the pathologist observed that UNI tended to focus on non-neoplastic areas containing smooth muscle and chronic inflammation, rather than tumor glands. In contrast, the pathologist noted that SPADE focused on areas with high grade dysplasia and invasive carcinoma. In one case SPADE also focused on regions with elongated nuclei (”spindle”-like). While spindle-cell morphology is not a canonical pattern of BRAF in colorectal cancer samples, spindle-cell patterns have been described in non-CRC BRAF-altered neoplasms such as melanoma [35]. In another case, the pathologist noted that UNI’s attention pattern appeared more random, focusing on necrotic regions rather than following the glandular architecture. Overall, in the six cases reviewed by our pathologist, UNI focused primarily on smooth muscle tissue and regions without tumor, whereas SPADE concentrated more on tumor regions.
Figure 4: Interpretability of SPADE on TCGA-CRC BRAF Task. Attention scores for a colorectal cancer WSI are visualized as a heatmap for the UNI and SPADE models. The deeper the red color, the higher attention the model puts on that region of the tissue. Sub-crops are selected to show the patterns with high attention at higher magnifications. A pathologist noted differences between UNI and SPADE heatmaps (see Section 4.4.6). Cancer outline delineated by a pathologist in black.
Next, we extracted the 50-highest attention patches from five different cases at 20x magnification for both SPADE and UNI (see Figure 5). We then provided patches to the pathologist, grouped by model and case, with model names obscured. Across the five cases, SPADE generally highlighted malignant glands, often with high-grade or cribriform architecture. UNI, in contrast, frequently highlighted lower-grade or fewer glands in its top 50 highest attention patches, and often emphasized stroma, fibrotic tissue, or inflammation rather than the tumor. In several cases, UNI identifies tiles with little or no carcinoma. The clinical impression suggests that SPADE is more tumor-focused, while UNI shifts attention to stromal or non-cancerous regions.
Figure 5: Highest Attention Patches SPADE vs UNI. We used the attention scores to extract the 50 highest-scoring attention patches from our model, SPADE, and compared them to the 50 highest-scoring attention patches from UNI. We then provided the resulting plot to a pathologist for review, with model names withheld. The pathologist noted differences between UNI and SPADE highest attention patches (see Section 4.4.6).
5 Discussion
Our study demonstrates that training on paired spatial transcriptomics and histology data substantially enhances the expressiveness of deep histology features across diverse cancer tasks. By leveraging gene expression data during contrastive pretraining, our framework improved performance on 17 out of 20 downstream tasks, including subtyping, survival, and biomarker prediction. These results highlight the value of multimodal supervision in building more robust histopathology foundation models.
Our method introduces a two-step clustering procedure coupled with a mixture-of-data-experts contrastive learning approach, enabling effective hard-negative mining for multi-organ histopathology representation learning. This training strategy enables learning a representation space that does not rely on trivial differences between histology tissues. Importantly, this approach can be extended to any histological dataset with multiple diseases and multiple organs to improve contrastive pretraining.